# DesCLIP: Robust Continual Adaptation via General Attribute Descriptions for Pretrained Vision-Language Models

### v1

@@ IV-B Comparison with State-of-the-art Methods
TABLE III: Ablation study of each component on Split CUB-200 [30].
@@
$\Delta$Last $\Delta$Avg $\Delta$Last $\Delta$Avg
@@
TABLE IV: Ablation study of different parameter selections on ImageNet-Subset [24] and CUB-200 [30].
$\lambda_{\textbf{IM}}$ threshold param $\gamma$ ImageNet-Sub [24] CUB-200 [30]
0 0.015 0.03 0.05 Last Avg Last Avg
0 82.8 86.0 - -
1.0 83.1 86.7 - -
2.0 82.9 86.4 - -
2.0 83.5 86.9 - -
2.0 84.2 87.3 76.9 82.0
2.0 83.8 87.1 - -
4.0 82.5 84.9 - -
0 - - 75.4 81.5
10.0 - - 77.8 82.8
15.0 - - 77.6 82.0
15.0 81.8 85.2 78.4 83.8
15.0 - - 78.0 83.5
20.0 - - 78.2 83.3
TABLE V: Results of different types of $\Psi(\mathbf{w},\mathbf{s})$ on Split CUB-200 [30]. $\varphi(\mathbf{w})$ represents using class-wise adapters [62] on rudimentary text embeddings.
$\Psi(\mathbf{w},\mathbf{s})$ $T=10$ $T=20$
@@
$\mathbf{w}$ 64.5 69.9 62.9 68.2
$\mathbf{w}+\alpha\cdot\varphi(\mathbf{w})$ 68.9 74.2 66.7 72.5
$\mathbf{w}+\alpha\cdot\mathbf{s}$ 77.5 82.9 74.8 81.0
$\mathbf{w}/\|\mathbf{w}\|+\alpha\cdot\varphi(\mathbf{w}/\|\mathbf{w}\|)$ 70.7 75.8 68.1 73.2
$\mathbf{w}/\|\mathbf{w}\|+\alpha\cdot\mathbf{s}$ 78.4 83.8 75.3 81.7
The accuracy results for continual learning, including ‘Last’ and ‘Avg’, are summarized in Table I. These results are derived from comprehensive experiments conducted on multiple datasets under varying incremental task settings. Additionally, Fig. 5 depicts the forgetting curves, comparing recent methods and providing a detailed evaluation of accuracy after each incremental task.
1) Performance on Coarse Dataset: CIFAR100 [29] is selected as the coarse dataset, where VLMs generally perform well on common classes but may struggle with blurred images. As shown in Table I, our method consistently achieves superior performance across all incremental task stages, obtaining an average accuracy (‘Avg’) of 90.1% under the setting of $T=10$. Notably, our approach surpasses ContinualCLIP [58] by a significant margin of 11.6% on the ‘Last’ metric. Moreover, it outperforms TaskRes-CL [5], MoE-Adapter [18], and SPU [10] by +3.3%, +2.4% and +1.6%, respectively.
2) Performance on Fine-grained Datasets: Table I demonstrate that our method consistently achieves the best performance across all incremental task stages on fine-grained datasets ImageNet-Subset [24] and CUB-200 [30]. On ImageNet-Subset, our method outperforms TaskRes-CL [5], SPU [10], and AttriCLIP [16] by +1.7%, +1.8%, and +5.7%, respectively. For CUB-200, our approach achieves the highest task accuracy at every stage. With $T=10$, our method surpasses SLCA by +4.5%, SPU by +2.2%, and TaskRes-CL by +3.4% on the ‘Last’ metric.
[…]
Figure 9: Visualizations of compliance and noncompliance attribute descriptions decided by AEF.
Figure 10: Visualizations of closest descriptions.
4) Zero-shot Degradation or Improvement? We propose a metric to evaluate the relative accuracy changes for all classes with respect to the zero-shot CLIP model after incremental tasks. Specifically, we calculate the class-wise accuracy difference between the zero-shot CLIP and the trained CLIP model after half of the incremental tasks (50 classes), denoted as $\Delta_{50|(50,zs)}$, and present the results in Fig. 6 (Upper). Our method achieves the highest averaged $\Delta_{50|(50,zs)}$, reaching +9.3. This demonstrates that our approach experiences the least zero-shot degradation while achieving superior performance, attributed to the more robust visual-textual relationships established during continual learning compared to the original vision-class text relationships.
5) Few-shot Performance: Fig. 7 illustrates the continual learning performance in few-shot scenarios on the CUB-200 dataset [30]. When the number of training samples per task is reduced from ‘full’ to ‘1/5’, our method maintains a competitive advantage, with an average performance drop of only -6.4. This decline is significantly smaller compared to TaskRes-CL’s -7.3 and SLCA’s -18.0, showcasing the robustness of our approach under challenging conditions.
IV-C Ablation Study
1) Effectiveness of Each Component: We conduct detailed ablation studies on various components of our DesCLIP framework. As shown in Table III, the zero-shot model is used as the baseline, and the relative improvements or declines are reported. It is evident that a fully fine-tuned visual encoder (V.E.(full)) suffers from catastrophic forgetting. Without data replay, merely fine-tuning the visual encoder (V.E.) partially (tuning the first MLPs in each Transformer block) proves inadequate for continual learning [10]. Based on V.E., by integrating instance matching (IM) with visually filtered attribute description embeddings, we achieve a relative improvement of 2$\sim$3over the zero-shot baseline. This improvement highlights the effectiveness of IM in mitigating forgetting, which typically arises from overfitting specific downstream task classes. Furthermore, the impact of text embedding calibration (TEC) is notable. The combination of V.E.+TEC+IM+TA, leveraging the filtered attribute description embeddings, yields the best performance, showcasing the synergy between these components.
2) Parameter Selection: Table IV presents the 10-task performance under different parameter configurations. Our analysis reveals that the effectiveness of the instance matching strength, $\lambda_{\textbf{IM}}$, is highly sensitive to the specific task context. Stronger instance matching proves advantageous in scenarios requiring finer granularity, such as CUB-200 [30]. On the other hand, the filtering of attribute description embeddings plays a crucial role. Excessively conservative filtering (e.g., $\gamma=0$) or overly stringent filtering (e.g., $\gamma=0.05$) leads to noticeable performance degradation.
Figure 11: Highly-relevant and filtered-out visual instances.
Figure 12: t-SNE visualizations on CIFAR100 [29]. Dots: visual representations; Stars: text embeddings.
3) Types of Text Embedding Calibration: We demonstrate the comparisons of different types of TEC in Table V. “$\Psi(\mathbf{w},\mathbf{s})=\mathbf{w}/\|\mathbf{w}\|+\alpha\cdot\mathbf{s}$” is the optimal solution, which is an calibration of the rudimentary text embedding in cosine space, which can best approximate the alignment of the representative shared attribute description embedding.
4) Generated Attribute Descriptions: We analyze the impact of the generated description amount on the performance of continual learning, as shown in Fig. 8. For coarse CIFAR100 [29], we only employed the auxiliary prompt “Please maintain diversity between descriptions” to generate $n_{dsc}$ attribute descriptions for each class. Our findings indicated that the model performed optimally when $n_{dsc}$ is set to 30. In contrast, for the fine-grained CUB-200 [30], we have discovered that incorporating additional auxiliary prompts, “Describe a/an [CLS]’s attributes from its beak, eyes, body, belly, tail, wings, breast, etc.” results in improved performance (red star in Fig. 8) even with a smaller $n_{dsc}$. This approach enhances the quality of the descriptions, provided that the language assistant possesses sufficient knowledge of the object’s details.
IV-D Visualizations
1) Description of Compliance and Noncompliance: In Fig. 9, we present examples visual image along with attribute description sentences that match and do not match the attribute descriptions filtered by AEF. In particular, different instances on the left, also from ‘Cardinal’, select the similar attribute feature description “a stout, bright orange beak” Furthermore, the highly matched attribute “crimson red feathers that cover its entire body” filtered by the first instance is not represented in the second instance.
2) GA Descriptions Closest to Calibrated Text Embeddings: In Fig. 10, we present the descriptions that are most closely aligned with the calibrated text embeddings. For each class, we showcase three of the top five closest attribute descriptions across several classes. Notably, some common visual attribute features are reflected in these descriptions, such as “slender, pointed wings; streaked brown and white feathers; a clean white belly; and darker hues on the wings and tail” for the “Brown Creeper” category.
3) Filtered-out Visual Instances: As shown in Fig. 11, it can be observed that, compared to the retained visual instances, the filtered visual instances lack relevance to both general attributes and class-specific information.
4) t-SNE Visualization: Fig. 12 presents the t-SNE visualizations of CLIP’s visual representations and text embeddings of each class. We employed distribution alignment to transform text embeddings into the visual representation space, rather than using the original text embedding space. Intuitively, compared to Zero-shot, TaskRes-CL [5] adjusts text embeddings in downstream incremental tasks without optimizing visual representations; SPU [10] optimizes visual representations but struggles to align visual-text representations for unfamiliar classes. In contrast, our method achieves superior alignment of visual-text representations in incremental tasks.
V Limitations

### v2

@@ IV-B Comparison with State-of-the-art Methods
TABLE II: Comparison of various methods on CIFAR100, CUB-200, and ImageNet-R ($T=10$) with CLIP ViT-B/16. FR. represents using feature replay. AS. represents using additional structure.
CIFAR100 CUB-200 ImageNet-R
Method FR. AS. Last Avg C. Last Avg C. Last Avg C.
Seq FT 46.3 24.2 45.7 39.7 50.8 34.3
ContinualCLIP 68.3 63.6 55.1 63.6 72.0 63.6
L2P 64.5 72.9 41.8 66.3 75.5 43.8
SLCA 71.5 79.8 59.1 50.6 58.4 62.7 72.1 76.2 57.1
AttriCLIP 67.0 77.8 60.3 50.8 65.4 63.1 70.4 73.7 61.0
SPU 75.8 84.3 58.7 67.8 75.6 59.8 76.8 82.4 57.5
TaskRes-CL 75.6 83.2 63.6 67.1 73.8 63.6 77.4 81.5 63.6
MoE-Adapter 77.8 84.9 62.2 68.3 76.2 62.9 80.1 83.7 60.7
RAPF 78.5 85.1 54.0 72.4 81.0 52.6 76.1 81.5 54.1
ENGINE 72.3 82.1 52.8 65.5 74.2 52.0 76.0 79.3 54.6
ENGINE(w/ GDA) 75.6 83.7 52.8 72.4 79.5 52.0 76.7 79.7 54.6
DesCLIP (Ours) 79.1 85.7 62.0 72.0 78.6 62.5 79.9 84.2 60.4
Figure 5: Per-stage average accuracies on OOD datasets.
Figure 6: Degradation or improvement evaluation on ImageNet-Sub with 10 incremental tasks. Upper: accuracy differences relative to zero-shot VLM. Bottom: accuracy differences relative to VLM previously learned on the first-half classes.
The accuracy results for continual learning, including ‘Last’ and ‘Avg’, are summarized in Table I. These results are derived from comprehensive experiments conducted on multiple datasets under varying incremental task settings.
1) Performance on Coarse Dataset: CIFAR100 is selected as the coarse dataset, where VLMs generally perform well on common classes but may struggle with blurred images. As shown in Table I, our method consistently achieves superior performance across all incremental task stages, obtaining an ‘Avg’ of 90.1% under the setting of $T=10$. Notably, our approach surpasses ContinualCLIP by a significant margin of 11.6% on the ‘Last’ metric. Moreover, it outperforms TaskRes-CL, MoE-Adapter, SPU, and OSN by +3.3%, +2.4%, +1.6%, and 0.9%, respectively.
2) Performance on Fine-grained Datasets: Table I demonstrates that our method consistently achieves the best performance across all incremental task stages on fine-grained datasets ImageNet-Sub and CUB-200. On ImageNet-Sub, our method outperforms TaskRes-CL [7], SPU [2], and AttriCLIP [16] by +1.7%, +1.8%, and +5.7%, respectively. For CUB-200, our approach achieves the highest task accuracy at every stage. With $T=10$, our method surpasses SLCA by +4.5%, SPU by +2.2%, and TaskRes-CL by +3.4% on the ‘Last’ metric.
3) Performance on OOD Datasets: Fig. 5 presents the per-stage performance comparison on OOD datasets. On FGVC-Aircraft, we evaluate two settings: Shuffle, where the task class order is randomly shuffled, and Hard, where the CL model is trained and evaluated following the original class order defined in FGVC-Aircraft. Under the Hard setting, the model is required to classify visually similar but previously unseen categories within a new task, such as distinguishing between Boeing 727-200 and Boeing 737-200. As shown in Fig. 5, our method consistently outperforms almost all compared approaches. Although ‘ENGINE w/ GDA’ achieves performance comparable to ours, it relies on stored distributional information to perform Gaussian Discriminant Analysis for constructing a visual classifier’s weights and bias, whereas our method does not store any information from previous tasks. In the Hard setting, our improvement is more pronounced, which gains +2.4% ‘LAST’ improvements compared to ‘ENGINE w/ GDA’. For other methods, fitting “unfamiliar” visual-text representations is even more detrimental to forgetting. In addition, we also report results on SKIN-40. Under both $T=5$ and $T=10$ settings, our method continues to deliver superior performance, while we observe that methods such as ENGINE and AttriCLIP almost forget all previously learned knowledge, with their ‘LAST’ scores close to or even lower than the zero-shot baseline of ContinualCLIP. These results demonstrate the OOD generalization capability of our method. It should be emphasized that this performance gain relies on proper GA descriptions to bridge the matching gap between visual and textual modalities, which constitutes the prerequisite for the effectiveness of our proposed AEF mechanism.
Figure 7: Loss variation of each steps on validation set of ImageNet-1K.
TABLE III: Ablation study of each component on CUB-200.
@@
$\DeltaLast Δ\DeltaAvg$ $\DeltaLast Δ\DeltaAvg$
@@
TABLE IV: Settings of threshold parameters $\delta_{d}$ and $\gamma$. The estimated values are highlighted by  .
Dataset CLIP ViT-L/14 CLIP ViT-B/16
$\delta_{d}$ $\gamma(e^{-2})$ $\delta_{d}$ $\gamma(e^{-2})$
CIFAR100 0.2 0.19 1.5 1.3 0.2 0.21 1.5 1.7
ImageNet-Sub 0.25 0.21 3 2.8
CUB-200 0.25 0.24 1.5 1.7 0.3 0.30 1.5 1.8
ImageNet-R 0.2 0.19 1.5 1.8 0.25 0.25 3 2.2
FGVC-Aircraft 0.2 0.22 3 2.0 0.25 0.27 3 2.0
SKIN-40 0.2 0.22 3 2.2 0.25 0.26 3 2.4
[…]
5) Zero-shot Degradation or Improvement? We propose a metric to evaluate the relative accuracy changes for all classes with respect to the zero-shot CLIP model after incremental tasks. Specifically, we calculate the class-wise accuracy difference between the zero-shot CLIP and the trained CLIP model after half of the incremental tasks (50 classes), denoted as $\Delta_{50|(50,zs)}$, and present the results in Fig. 6 (Upper). Our method achieves the highest averaged $\Delta_{50|(50,zs)}$, reaching +9.3. This demonstrates that our approach experiences the least zero-shot degradation while achieving superior performance, attributed to the more robust visual-textual relationships established during continual learning compared to the original vision-class text relationships. As shown in Fig. 7, during training, our method maintains a lower validation loss on the balanced subset of ImageNet-1K compared to the standard slow leaner (SL) [1], thereby mitigating the risk of forgetting pretrained knowledge.
Figure 8: Performance comparison in few-shot case.
Figure 9: Hyperparameter estimation and cross-dataset evaluations.
Figure 10: Ablation study of generated GA description amount.
6) Few-shot Performance: Fig. 8 illustrates the continual learning performance in few-shot scenarios on the CUB-200 dataset. When the number of training samples per task is reduced from ‘full’ to ‘1/5’, our method maintains a competitive advantage, with an average performance drop of only -6.4. This decline is significantly smaller compared to TaskRes-CL’s -7.3 and SLCA’s -18.0, showcasing the robustness of our approach under challenging conditions.
IV-C Ablation Study
[…]
2) Hyperparameter Analysis: The hyperparameter setting is crucial for the effectiveness of the proposed AEF mechanism. Given the inherent differences in visual–textual similarity across downstream tasks, it is critical to efficiently determine suitable hyperparameters for CL. In our study, we first obtain preliminary estimates of the threshold parameters $\delta_{d}$ and $\gamma$, and then refine their values through empirical tuning to ensure better performance across tasks. Specifically, the estimation process is as illustrated in Fig. 9(a). In the first CL task, we compute the similarity distribution between training images and their corresponding class-text embeddings, and take the 5th percentile (P05) of this distribution as the estimated value $\hat{\delta_{d}}$. Meanwhile, for each visual instance, we compute the similarity with its most aligned GA description embedding, $\mathcal{H}_{i}=\left\langle\frac{\mathbf{z}_{i}}{\|\mathbf{z}_{i}\|},\frac{\mathbf{h}_{i}}{\|\mathbf{h}_{i}\|}\right\rangle,$ and estimate $\hat{\gamma}$ via $\hat{\gamma}=\mathbb{E}_{i=1:N_{t=1}}\left(\mathcal{H}_{i}-CS_{i}\right)\cdot 0.7,$ which provides a reasonable initialization and is further adjusted in practice to mitigate the effect of noise on the AEF mechanism in downstream tasks. The downstream-task settings of specific threshold parameters and estimated parameters (highlighted) are reported in TABLE. IV.
We analyze the robustness of hyperparameters obtained through this unified estimation strategy and compare the results with those trained under dataset-specific hyperparameters. In addition, we provide further comparisons to show that, even with a unified hyperparameter setting strategy, our method still surpasses competing approaches. The CL performance across multiple datasets is illustrated in Fig. 9(b). In particular, we evaluate the specific configuration for each dataset (#1), the estimated threshold hyperparameters (#2–3), different fixed hyperparameters (#4–10), and two recent CL methods (#11–12). These results demonstrate that our approach exhibits low dependence to hyperparameter selection, since both estimated and fixed settings still achieve performance that is almost consistently superior to recent CL baselines. Moreover, we observe that fine-grained downstream tasks, such as CUB-200 and FGVC-Aircraft, generally favor higher values of $\lambda_{\textbf{IM}}$. However, even with a fixed loss coefficient $\lambda_{\textbf{IM}}$, competitive CL performance can still be achieved. On the other hand, filtering of GA description embeddings proves to be a key factor: overly conservative filtering (e.g., $\gamma=0$) or excessively strict filtering (e.g., $\gamma=0.05$) leads to a noticeable drop in performance.
Figure 11: Visualizations of Left: compliance and noncompliance GA descriptions decided by Eq. 12; Right: Filtered-out visual instances by Eq. 10
3) Generated Attribute Descriptions: We analyze the impact of the generated description amount on the performance of continual learning, as shown in Fig. 10. For coarse CIFAR100, we only employed the auxiliary prompt “Please maintain diversity between descriptions” to generate $n_{dsc}$ attribute descriptions for each class. We find that the model performed optimally when $n_{dsc}$ is set to 30. In contrast, for the fine-grained CUB-200, we have discovered that incorporating additional auxiliary prompts, “Describe a/an <class name>’s attributes from its beak, eyes, body, belly, tail, wings, breast, etc.” results in improved performance (red star in Fig. 10) even with a smaller $n_{dsc}$. This approach enhances the quality of the descriptions, provided that the language assistant possesses sufficient knowledge of the object’s details.
Moreover, redundancy was found among the generated GA descriptions. In CIFAR100, for instance, only a limited number of GA embedding candidates were selected as highly relevant to the corresponding visual instances for each class. To further optimize the GA candidate pool, we explored a static pre-filtering strategy before training. Specifically, we applied a fixed AEF-based selection to evaluate the matching confidence of each GA embedding without updating the CLIP model. For each class, GA candidates were ranked by their selection frequency, and the top-$Q$ candidates were retained to form a reduced GA pool. As shown in Fig. 10, this pre-filtering achieves a modest yet consistent improvement in CL performance, particularly on CIFAR100. Notably, using only 10 pre-filtered GA descriptions surpasses the performance achieved with 20 unfiltered candidates. In contrast, on CUB-200, where GA descriptions are inherently more distinctive and less redundant, the impact of pre-filtering is less significant. These results suggest that while GA diversity is crucial, a compact set of high-confidence descriptions is sufficient to guide instance matching effectively.
Figure 12: t-SNE visualizations on CIFAR100. Dots: visual representations; Stars: text embeddings.
IV-D Visualizations
1) Description of Compliance and Noncompliance: In Fig. 11(Left), we present visual instances alongside GA descriptions that either match or do not match the attribute features filtered by AEF, together with their corresponding similarity scores. This illustrates the degree of visual-text alignment for each GA under both compliant and non-compliant conditions. Compared to manually designed prompts, the GA descriptions filtered by AEF exhibit significantly stronger visual relevance, especially in specialized domains such as dermatological diagnosis.
2) Filtered-out Visual Instances: As shown in Fig. 11(Right), it can be observed that, compared to the retained visual instances, the filtered visual instances lack relevance to both general attributes and class-specific information.
3) t-SNE Visualization: Fig. 12 presents the t-SNE visualizations of CLIP’s visual representations and text embeddings of each class. We employed distribution alignment to transform text embeddings into the visual representation space, rather than using the original text embedding space. Intuitively, compared to Zero-shot, TaskRes-CL [7] adjusts text embeddings in downstream incremental tasks without optimizing visual representations; SPU [2] optimizes visual representations but struggles to align visual-text representations for unfamiliar classes. In contrast, our method achieves superior alignment of visual-text representations in incremental tasks.
TABLE V: Comparison of different description sources and their CL performances on CUB-200.
Description Types LLM Types Num. / Class Avg (%)
CUB-attributes1 [74] - 300 67.8
CUB-attributes2 [74] - 30 65.5
DRPs GPT-3 30 81.0
DRPs GPT-4 30 82.9
DRPs DeepSeek-R1 30 82.2
DRPs w/ auxiliary prompts DeepSeek-R1 30 83.2
DRPs w/ auxiliary prompts GPT-4 30 83.8
Figure 13: Cross-task validations on Image Retrieval and Image Caption with CLIP ViT-B/16.
V Discussions
V-A In-depth Study of Text Generation Strategies
1) GA Description Source: We conducted an in-depth study on the impact of GA descriptions from different sources on continual learning performance using the CUB-200 dataset. As shown in Table V, we compared several types of descriptions: CUB-attributes curated by Reed et al. [74], GA descriptions generated by GPT-3 [75], DeepSeek-R1 [76], and GPT-4 [54]. Our results indicate that the CUB-attributes, which focus exclusively on object-level attributes and omit explicit class identifiers, result in inferior performance. This outcome suggests that CLIP also benefits from class-level semantic cues, which are absent in such formulations. In contrast, GA descriptions produced by other LLMs, such as DeepSeek-R1 [76], achieve comparable continual learning performance, especially when auxiliary prompts are employed to guide the model in emphasizing key visual attributes.
2) Comparison of LLM-based Description Generation Method: Table VI reports the performance gains obtained by replacing handcrafted templates with LLM-generated GA descriptions. Zero-shot CLIP with handcrafted prompts serves as the baseline. We compare several representative LLM-based text generation methods, including CuPL [26], DCLIP [27], and advanced strategies such as FuDD [29] and MPVR [31], with the following variants:
  • FuDDAS: Differential descriptions generated using all seen classes.
  • FuDDCL: Differential descriptions generated only from classes observed in the current task.
  • MPVRG: Rich semantic prompts generated via meta-prompting using GPT-3.5.
  • MPVRM: Rich semantic prompts generated via meta-prompting using Mixtral-7B [77].
We observe that replacing handcrafted templates with these LLM-based textual descriptions consistently improves performance, both in zero-shot settings and under CL methods. However, all variants remain inferior to our proposed DesCLIP. Furthermore, Table VII presents results where the generated texts from these representative strategies are utilized to construct the GA description candidate pool, essentially functioning as variants of DesCLIP. Based on our observations, advanced LLM-based methods like MPVR tend to produce more detailed and comprehensive semantic descriptions, while FuDD emphasizes generating discriminative attributes to resolve inter-class ambiguities. Compared to simpler generation methods such as CuPL and DCLIP, these advanced approaches yield significantly better performance in CL scenarios.
TABLE VI: ‘Last’ (T=10T{=}10) improvement compared to zero-shot CLIP ($\%$). ‘Handcraft’ represents using the template ‘A photo of a <classname><class~name>.’ to obtain class-text embeddings. The values in parentheses indicate the relative gain compared to using handcraft template.
Method CL Training ImageNet-Sub CUB-200
LLM-generated
texts
CuPL +1.1 +0.8
DCLIP +3.8 +1.2
FuDDCL +1.7 +1.6
FuDDAS +4.1 +3.6
MPVRG +3.6 +3.0
MPVRM +3.3 +2.5
CL methods
+ Handcraft
AttriCLIP +3.9 -8.4
TaskRes-CL +7.9 +14.1
SPU +7.8 +15.3
Combined TaskRes-CL + CuPL +7.9 (+0.0) +14.5 (+0.4)
TaskRes-CL + FuDDCL +8.5 (+0.6) +15.9 (+1.8)
TaskRes-CL + MPVRG +7.5 (-0.4) +15.3 (+1.2)
SPU + CuPL +7.1 (-0.7) +15.8 (+0.5)
SPU + FuDDCL +8.0 (+0.2) +16.3 (+1.0)
SPU + MPVRG +7.1 (-0.7) +16.1 (+0.8)
DesCLIP +9.6 +17.5
TABLE VII: Comparison of different GA description candidate pool construction methods and pool sizes. We report ‘Avg’ and ‘Last’ ($T=10$). ‘DRP’ represents using additional auxiliary prompts.
GA candidate pool
Construction
Pool size ImageNet-Sub CUB-200
@@ V-A In-depth Study of Text Generation Strategies
+ DRP 10 82.7 86.1 77.9 82.7
+ DRP 30 84.2 87.3 77.6 82.5
+ DRP 30 79.0 83.3
+ DRP 20 80.1 84.2
+ CuPL 30 82.2 84.9 76.1 80.3
+ DCLIP 30 83.8 87.0 78.5 82.9
+ FuDDCL 85.4 88.6 77.1 81.2
+ MPVRG 30 83.4 85.8 77.9 83.0
+ MPVRG 100 85.2 88.1 79.9 83.2
V-B Preservation of Cross-Task General Knowledge
Pre-trained VLMs are applicable not only to image recognition but also to a broad spectrum of vision–language tasks. Therefore, under the CL setting, it is crucial to examine whether cross-task general knowledge can be effectively preserved. To this end, we evaluate performance drift on two representative downstream tasks. For image retrieval, we utilize the CL-adapted CLIP weights and measure both image-to-text (I2T) and text-to-image (T2I) retrieval performance on the MS-COCO [78] and Flickr30K [79] benchmarks. For image captioning, we adopt the ClipCap [80] framework by reusing the continually adapted CLIP weights and assess caption quality on the COCO-Captions [81] dataset, reporting BLEU-4 [82] and CIDEr [83] scores. As illustrated in Fig.13, DesCLIP not only achieves superior performance on the CL tasks, but also maintains competitive performance on the image retrieval task. Compared to SL and SPU, it exhibits a smaller drop in both R@1 and R@5 scores, remaining close to the original zero-shot CLIP performance. Similarly, DesCLIP demonstrates the least degradation in image captioning quality, indicating its strong ability to preserve task-agnostic general knowledge throughout the continual learning process.
V-C Limitations